The Pointer Appliance Company is investigating the additional of a new and improved pulsating blender to its line of consumer appliances.   The product chops, grinds, grates and blends smoothies twice as fast as all other blenders on the market. It has a 10 year warranty on the motor and offers a replacement 5 cup glass blender cup for $10.99. The blender is going to require additional expenditures for the production line on the part of the Pointer Company. Management wants to determine what to price the blender at. The total fixed cost associated with manufacturing the blenders is $525,000. The variable cost is $17.99 per unit. Given the costs for adding the blender, management is considering selling the blender at one of these prices: $25.99, $34.99 and $49.99. Calculate the breakeven point in units at each of the selling prices: $25.99, $34.99 and $49.99 and show the calculations. Through your research you know that this is a tough competitive market with many competitors. Most competitors have between 2-8% of the market and the market appears to be growing at a 5% rate per year. The total market for blenders of this type was 1,500,000 units last year. Given that the market leader is selling its blender for $39.99, which price would you recommend and why?
Answer #1

If the break-even number of units is X then X = fixed cost/(selling price per unit - variable cos per unit)

a). Selling price = 25.99

X = 525,000/(25.99-17.99) = 65,625

b). Selling price = 34.99

X = 525,000/(34.99-17.99) = 30,882.35

c). Selling price = 49.99

X = 525,000/(49.99-17.99) = 16,406.25

The market leader would be the price setter for the appliance, so the price has to be lower than the price at which it sells but not so low that profits depend significantly upon volume (as it is a new product), so the appropriate price is $34.99.